# Diagnosing Vision Language Models’ Perception by Leveraging Human Methods for Color Vision Deficiencies

### v1

@@ Abstract
Large-scale Vision Language Models (LVLMs) are increasingly being applied to a wide range of real-world multimodal applications, involving complex visual and linguistic reasoning. As these models become more integrated into practical use, they are expected to handle complex aspects of human interaction. Among these, color perception is a fundamental yet highly variable aspect of visual understanding. It differs across individuals due to biological factors such as Color Vision Deficiencies (CVDs), as well as differences in culture and language. Despite its importance, perceptual diversity has received limited attention. In our study, we evaluate LVLMs’ ability to account for individual level perceptual variation using the Ishihara Test, a widely used method for detecting CVDs. Our results show that LVLMs can explain CVDs in natural language, but they cannot simulate how people with CVDs perceive color in image based tasks. These findings highlight the need for multimodal systems that can account for color perceptual diversity and support broader discussions on perceptual inclusiveness and fairness in multimodal AI.
@@
Kazuki Hayashi        Shintaro Ozaki Yusuke Sakai        Hidetaka Kamigaito       Taro Watanabe Nara Institute of Science and Technology (NAIST), Japan {hayashi.kazuki.hl4, ozaki.shintaro.ou6 sakai.yusuke.sr9, kamigaito.h, taro.watanabe}@naist.ac.jp
@@ 1 Introduction
Large-scale Vision Language Models (LVLMs) Liu et al. (2024); Abdin et al. (2024); Ye et al. (2024); Bai et al. (2025) possess advanced reasoning capabilities that jointly process images and text, and are being integrated into real-world applications Zitkovich et al. (2023); Hao et al. (2024); Morita et al. (2024). These models should be designed with consideration for perceptual variations, ensuring accessibility and usability for individuals with differing perceptual characteristics. Color is a fundamental dimension of visual perception and a frequent reference point in everyday communication and decision making Stroop (1935). However, color perception varies across individuals due to biological and cultural differences. Color vision deficiencies (CVDs) are a common example, making it difficult to distinguish colors such as red and green Roberson and Hanley (2007); Birch (2012b).
Figure 1: A communication failure caused by color vision deficiency, where the model’s color-based directions may not be understood by some users.
Current LVLMs are trained mainly on image and text corpora that assume normal color vision Paik et al. (2021); Samin et al. (2025); Rahmanzadehgervi et al. (2024), potentially overlooking the needs of users with diverse perceptual profiles. As Figure 1 illustrates, an LVLM based navigation assistant that instructs a traveller to “Take the green line” may leave a red and green color blind user unable to act. Such mismatches degrade user experience and raise concerns about fairness and inclusivity in real-world deployments Kawakita et al. (2024).
@@
Table 1: Ground-truth answers for each Ishihara plate under different vision conditions, based on clinical interpretations of the Ishihara color test Ishihara (1917).
@@
What is Ishihara Test? … visual test used to detect color-vision deficiencies, especially red–green types such as protanopia and deuteranopia. It presents a pattern of red and green dots forming a hidden number; individuals with normal vision can read the number, whereas those with certain CVDs cannot.
What is Protanopia? color blindness that affects a person’s ability to distinguish between red and green colors. In people with protanopia, the cone cells in the retina that are responsible for perceiving red and green colors are not functioning properly.
Table 2: Representative excerpts from mPLUG-Owl3 showing detailed knowledge of the Ishihara test and CVDs. The full outputs are available in Appendix C.
@@
Base You are Protanopic.
@@
Linguistic Support You are Protanopic.
@@
Visual Support You are Protanopic.
@@
Table 3: Examples of protanopia prompts. Other CVD types are provided in Appendix B.
To address this, we analyze how well LVLMs handle diverse human color perception. We focus on the Ishihara Test Ishihara (1917), a widely used and trusted diagnostic tool for assessing CVDs, and use it to diagnose the visual understanding of LVLMs under various CVD scenarios. Our experimental analysis shows that although LVLMs possess factual knowledge of CVDs, they fail to simulate and understand how people with color vision deficiencies perceive colors. This finding reveals a crucial limitation of current LVLMs in aligning with diverse human perceptual experiences and highlights the need for developing models which more fully capture diverse color perceptions.
@@
Table 4: Accuracy (%) of each LVLM under simulated Vision types: Normal and three CVD types, with support settings: Base, Ling. (Linguistic), and Vis. (Visual). Vis. was used only for models supporting multi-image input.
2 Preliminaries
Ishihara Test
is a widely used standard tool for diagnosing red and green CVDs Ishihara (1917). It uses circular plates commonly called Ishihara plates, composed of colored dots that form numerals which appear differently depending on the viewer’s color vision type. For instance, the middle dot plates in Table 1 shows “8” under normal color vision but “3” under red and green CVDs.
Color Vision Deficiency Types Birch (2012a):
Protanopia : Reds appear darker, and it is difficult to distinguish between red, green, and brown. This is due to the absence of L-cones.
Deuteranopia : Greens may appear beige, causing red and green confusion like Protanopia but without dark reds. Caused by M-cone loss.
Tritanopia : Blues shift toward green, and yellows may appear pink or gray, making blue–yellow distinctions challenging. This is linked to the absence of S-cones.
Pilot Study.
Before evaluating whether LVLMs simulate human color perception, we first confirmed that they understand basic concepts of CVDs and the Ishihara Test. We used two prompts: (i) show an Ishihara plate and ask “What is this image used for?”, and (ii) ask for definitions such as “Please explain what Protanopia is as a type of color blindness.” As shown in Table 2, the models responded correctly, confirming readiness for the main tasks. Full outputs are in Appendix C.
Figure 2: Violin plots of mPLUG-Owl3 perplexity on Ishihara digit prediction across four vision types and three prompt settings. Lower perplexity (from force-decoding the gold answer) indicates higher model confidence.
3 Experiments and Analysis
3.1 Research Questions
We use the Ishihara Test as with humans to analyze how LVLMs perceive and simulate color vision across different task settings [^1] We use 25 out of 38 plates that require numeral responses to ensure consistent analysis across visual conditions.. Our goal is to systematically investigate LVLMs’ color perception across three analytical levels: output behavior, confidence, and internal representation.
RQ1: At the generation level, can LVLMs generate digit responses that resemble those typically reported by people with color-vision deficiencies on Ishihara plates? (§3.3)
RQ2: At the confidence level, how accurate are the model’s outputs across vision conditions, and how confident (via perplexity) is the model when producing them? (§3.4)
RQ3: At the internal representation level, what layerwise probability patterns emerge when LVLMs infer CVD types from Ishihara responses? (§3.5)
3.2 Settings
LVLMs.
We evaluate the following LVLMs: LLaVA-NeXT Liu et al. (2024), Llama-3.2 Grattafiori et al. (2024), mPLUG-Owl3 Ye et al. (2024), Phi-3.5 Abdin et al. (2024), Qwen2.5-VL Bai et al. (2025), and GPT-4o OpenAI et al. (2024). GPT-4o represents a proprietary model, and the others are chosen for their knowledge of CVDs and the Ishihara Test in pilot study of § 2. Appendix A provides more details.
Prompts.
We evaluate LVLMs using Ishihara plates and three prompt types, as shown in Table 3. The Base prompt states only the condition. The Linguistic Support prompt adds a brief explanation of how the condition affects color perception. The Visual Support prompt presents an example image showing how the plate appears to individuals with the condition, before the question.

### v2

@@ Abstract
Large-scale Vision-Language Models (LVLMs) are being deployed in real-world settings that require visual inference. As capabilities improve, applications in navigation, education, and accessibility are becoming practical. These settings require accommodation of perceptual variation rather than assuming a uniform visual experience. Color perception illustrates this requirement: it is central to visual understanding yet varies across individuals due to Color Vision Deficiencies, an aspect largely ignored in multimodal AI. In this work, we examine whether LVLMs can account for variation in color perception using the Ishihara Test. We evaluate model behavior through generation, confidence, and internal representation, using Ishihara plates as controlled stimuli that expose perceptual differences. Although models possess factual knowledge about color vision deficiencies and can describe the test, they fail to reproduce the perceptual outcomes experienced by affected individuals and instead default to normative color perception. These results indicate that current systems lack mechanisms for representing alternative perceptual experiences, raising concerns for accessibility and inclusive deployment in multimodal settings.
[…]
@@
Kazuki Hayashi    Shintaro Ozaki Yusuke Sakai    Hidetaka Kamigaito    Taro Watanabe Nara Institute of Science and Technology (NAIST), Japan hayashi.kazuki.hl4@is.naist.jp ozaki.shintaro.ou6@naist.ac.jp {sakai.yusuke.sr9, kamigaito.h, taro.watanabe}@is.naist.jp
@@ 1 Introduction
Large-scale Vision-Language Models (LVLMs) liu2024llavanext; abdin2024phi3technicalreporthighly; ye2024mplugowl3longimagesequenceunderstanding; bai2025qwen25vltechnicalreport extend language models with visual perception and enable unified multimodal reasoning. As LVLMs move from research prototypes to real-world interactive systems, they are increasingly deployed in navigation, education, and assistive technologies pmlr-v229-zitkovich23a; jimaging10050103; morita2024contextawaresupportcolorvision; zhou2025autovla. In such applications, LVLMs serve as a medium through which users access and interpret visual information in their environment.
Figure 1: An example of communication failure due to CVDs: a user with CVD may not correctly interpret the model’s color-based instructions. This highlights the need for LVLMs to account for perceptual differences in real-world interactions.
A key yet underexplored assumption in such deployments is that users share a similar visual experience of the world. In reality, visual perception varies across individuals. Color vision in particular shows substantial variability due to biological factors Emery2019-hi, and Color Vision Deficiencies (CVDs) are a common form of such diversity, impairing discrimination between colors such as red and green Birch2012-xg; roberson2007color. These differences are not only perceptual but also functional: color vision affects wayfinding, diagram and chart reading, medical interpretation, and user interface navigation. As illustrated in Figure 1, an LVLM-based navigation assistant instructing a traveler to “Take the green line” may fail to provide actionable information for a red–green color blind user, raising concerns about inclusion, accessibility, and fairness in real deployments Kawakita2024. Systems designed around a “standard” visual experience may silently exclude users with different perceptual characteristics, even when linguistic content remains fully accessible.
Current LVLMs are trained largely on web-scale image-text corpora that implicitly assume typical color vision paik-etal-2021-world; samin-etal-2025-colorfoil; Rahmanzadehgervi_2024_ACCV. As a result, these models acquire substantial knowledge about color, color blindness, and clinical diagnostic tools. However, whether such knowledge enables the model to simulate alternative visual percepts remains unclear. This distinction is central: explaining what CVD is in language is not the same as perceiving colors as a CVD user would. To our knowledge, the ability of LVLMs to reproduce perceptual diversity has not been systematically investigated.
To address this, we leverage the Ishihara Test ishihara1917test, a clinically validated tool widely used for diagnosing CVDs in humans. Rather than using Ishihara solely as a screening instrument, we repurpose it as a diagnostic probe for LVLMs. Because Ishihara plates are designed to elicit different perceptual responses across color vision types, they allow direct comparison between model outputs and established human perceptual categories. In addition, the test provides a standardized visual stimulus that isolates color-based perceptual variation without cultural or linguistic interference, and its clinical validation over a century allows perceptual differences to be evaluated without new annotation pipelines or subjective rating procedures.
We evaluate LVLM behavior through three complementary perspectives. At the generation level, we assess whether the model outputs the digit perceived by users with different CVD types. At the confidence level, we analyze perplexity under simulated CVD conditions to determine whether uncertainty patterns align with human perceptual tendencies. At the representation level, we apply a layer-wise LogitLens probing nostalgebraist2020logitlens method to assess whether internal activations encode distinctions between vision types. Together, these perspectives allow us to analyze how perceptual variation interacts with LVLMs’ generative, probabilistic, and latent processing mechanisms.
Our findings show that although LVLMs possess strong linguistic knowledge about CVDs and can describe the Ishihara Test, they fail to reproduce the perceptual outcomes experienced by individuals with CVD. Even when given additional linguistic descriptions or visual exemplars, the models default to normative color perception, indicating insufficient alignment between linguistic knowledge and perceptual grounding. These results suggest that current LVLMs lack mechanisms for representing alternative perceptual experiences, which has implications for accessibility, fairness, and human-centered deployment in multimodal AI as such systems increasingly mediate visual information.
2 Background
2.1 Color Vision Deficiency and Ishihara Test
Color vision in humans relies on three types of cone photoreceptors: red, green, and blue, each sensitive to long, medium, or short wavelengths. CVDs affect about 8% of males and 0.5% of females Birch2012-xg. When one type of cone is missing or doesn’t work properly, large parts of the color spectrum lose their typical hue or brightness, resulting in CVDs. The three main types are summarized below:
Protanopia
Absence of long wavelength (L) cones makes reds appear very dark and compresses the differences among many reds, greens, and browns; purples are often mistaken for blue.
Deuteranopia
Absence of medium wavelength (M) cones leaves brightness almost unaffected but shifts greens toward beige, producing strong red and green confusion.
Tritanopia
Absence of short wavelength (S) cones narrows the blue to yellow axis; blues drift toward green, cyans toward gray, and yellows toward pink or light gray, greatly reducing blue versus yellow separability.
@@ 2.1 Color Vision Deficiency and Ishihara Test
Table 1: Ground-truth answers for each Ishihara plate under different vision conditions, based on clinical interpretations of the Ishihara color test ishihara1917test.
To screen for CVDs, the Ishihara Test (ishihara1917test) has been widely used for decades in clinical and research settings Fanlo_Zarazaga2019-wc. Each plate shows a figure made of colored dots. People with normal vision usually identify it, while those with CVDs may see a different number or nothing. For instance, the plate in Table 1 looks like “8” to a person with normal vision, but is often read as “3” by people with red and green deficiencies Birch2012-xg. By using typical color confusions across CVDs, the Ishihara Test remains a fast and reliable tool for identifying these conditions.
2.2 CVDs and Ishihara Tests in Recent Research
Despite its limited size, the Ishihara Test remains a clinically validated and reliable tool for CVD detection. CHEN1995463 applied automated analysis to recognize Ishihara patterns, enabling automated color vision testing. Later work confirmed that digital versions retain reliability across devices zhao2019comparison; klinke2024comparative. computers14020034; grover2025huemanityprobingfinegrainedvisual introduced synthetic Ishihara-style datasets to assess model performance on digit or character recognition under red and green confusion. CVDs compensation methods also use Ishihara principles to enhance accessibility for people with color vision deficiency. Key approaches include physiologically based daltonization 5290741, eigenvector-based recoloring s19102250, and neural real-time Ishihara recoloring inbook. These works highlight Ishihara plates as diagnostic tools and foundations for assistive color algorithms.
@@ 2.2 CVDs and Ishihara Tests in Recent Research
[…]
Table 2: Excerpts from mPLUG-Owl3 showing knowledge of the Ishihara test and CVDs. The full outputs are available in Appendix B.
2.3 Color Understanding in LVLMs
Early VQA models balanced_vqa_v2 explored color-related questions but relied on priors over visual understanding, yielding inconsistent results. Agrawal_2018_CVPR; Kervadec_2021_CVPR proposed debiased datasets to reduce language bias and improve visual grounding. Transformer-based models pmlr-v139-radford21a; pmlr-v139-kim21k enhanced image-text alignment and performed well on some tasks yet still struggle with fine-grained color distinctions samin-etal-2025-colorfoil. Although recent LVLMs exhibit strong visual reasoning, their color understanding remains limited by insufficient supervision for subtle or ambiguous cases. burapacheep2024colorswapcolorwordorder; shahgir2024illusionvqachallengingopticalillusion found that models struggle with basic color-word grounding and are easily misled by perceptual illusions. liang2025colorbenchvlmsunderstandcolorful evaluated diverse color tasks and showed even large models struggle with shifts and ambiguity, revealing persistent limitations.
3 Experimental Setup
3.1 Pilot Study
Before analyzing our main question of whether LVLMs can simulate aspects of human color perception, we first verified that they possess basic language-level knowledge about color vision deficiencies and the Ishihara Test. To test this prerequisite, we used two types of prompts: (i) asking an Ishihara plate “What is this image used for?”, and (ii) requesting definitions such as “What is Protanopia as a type of color vision deficiency?”. As shown in Table 2, all evaluated models were judged to respond correctly. This confirms that they satisfy the necessary prerequisite knowledge for the analyses that follow. Appendix B provides full outputs.
3.2 Research Questions
We use the Ishihara Test as with humans to analyze how LVLMs simulate color vision across task settings [^1] We use 25 out of 38 plates that require numeral responses to ensure consistent analysis across visual conditions.. In our research, “color perception” means the model’s ability to respond like humans with specific CVDs, rather than identifying colors. We investigate this across three levels: generation, confidence, and internal representation.
@@ 3.2 Research Questions
Base {Image: Ishihara plate}
You are Protanopic.
@@
Linguistic Support {Image: Ishihara plate}
You are Protanopic.
@@
Visual Support {Reference Image: Ishihara plate}
You are Protanopic.
@@
Doctor-Style {Image: Ishihara plate}
[…]
Choose from the following options:
- Normal - Protanopia - Deuteranopia - Tritanopia Please output your answer.
Table 3: Examples of prompts used in our experiments. Base, Linguistic, and Visual prompts are used in RQ1-RQ2 for digit generation, and Doctor-Style is used in RQ3 for vision-type inference. All include an Ishihara plate as visual input.
RQ1 (Generation level).
[…]
RQ2 (Confidence level).
Do model confidence patterns vary across vision conditions, and do uncertainty signals (measured via perplexity) reflect condition-specific perceptual ambiguity observed in humans? This probes whether LVLMs assign different likelihoods to Ishihara responses depending on simulated vision type.
RQ3 (Representation level).
Do internal layerwise activation patterns encode distinctions consistent with color vision deficiency categories? This examines whether alternative perceptual states are represented at the latent level, even if not expressed in the final output.
3.3 Task Definition
The task is inspired by the human Ishihara Test: given an Ishihara plate, the model is instructed to simulate a target vision condition and report the perceived digit. We simulate four conditions: {Normal, Protanopia, Deuteranopia, Tritanopia}. Since the Ishihara Test does not assess Tritanopia, the correct digits for Normal and Tritanopia coincide. We exploit this property by treating Tritanopia as a control condition: if a model meaningfully incorporates the instruction “You are Tritanopic,” its output distribution or confidence should differ from Normal despite identical ground-truth digits. As shown later, most LVLMs instead return nearly identical responses for the two conditions, suggesting weak sensitivity to instructed perceptual states.
Prompts
To analyze whether LVLMs can simulate CVDs under different instructional prompts, we use Ishihara plates and design four types of prompts: Base, Linguistic Support, Visual Support and Doctor-Style, as shown in Table 3. We tested multiple variants within each prompt type but observed no substantial differences. Thus, we use one representative set in our main experiments and report the variant analysis in Appendix C.
Base Prompts
These provide only the condition, e.g., “You are Protanopic”, and ask the model to report the number in the Ishihara plate. This tests whether the model can simulate CVD perception specified by the condition.
Linguistic Support
In addition to the Base prompt, these add a short description of the impairment, e.g., “red tones appear darker,” to assess whether textual context aids prediction.
Visual Support
These include a brief condition description and a reference example showing what number appears in another plate, before asking the model to identify the number in a new image. This tests whether few-shot visual examples help simulate CVD perception.
Doctor-Style
Instead of generating digits, the model infers the vision type (Normal, Protanopia, Deuteranopia, or Tritanopia) from a given response. Used in RQ3, this prompt avoids tokenization issues with multi-digit numerals (e.g., “18” $\rightarrow$ “1”, “8”) by using diagnosis labels, which are consistently tokenized and clearly distinguishable across models.
@@ 3.3 Task Definition
Table 4: Digit Accuracy (%) of each LVLM under simulated Vision types: Normal and three CVD types, with support settings: Base, Ling. (Linguistic), and Vis. (Visual). Vis. was used only for models supporting multi-image input. Plates for which no correct digit exists (marked as “N/A”in the Gold answers in Table 5) were excluded from the accuracy calculation.
3.4 Evaluation Metrics
We quantify each research question using a dedicated and interpretable metric for each aspect.
Digit Accuracy (RQ1)
The proportion of Ishihara plates on which the model’s generated digit matches the ground‑truth numeral under the specified vision condition.
Per‑Token Perplexity (RQ2)
For each plate, we concatenate the image, prompt, and gold numeral answer into a single input sequence. All tokens preceding the answer are masked with ignore_index, ensuring that only the answer tokens contribute to the loss. The answer consists of $L$ tokens, denoted as $\boldsymbol{x}_{1:L}$. We compute their token‑level log‑probabilities
$v_{i}=\log P\left(x_{i}\mid\textit{prompt},\,\boldsymbol{x}_{1:i-1}\right),$
and the average negative log‑likelihood
$\mathcal{L}\;=\;-\frac{1}{L}\sum_{i=1}^{L}v_{i},\quad\mathrm{PPL}\;=\;\exp(\mathcal{L}).$
A lower $\mathrm{PPL}$ indicates higher confidence in the forced-decoded answer.
Layer‑Wise Diagnosis Probability (RQ3)
This approach builds on prior work showing that intermediate representations influence model predictions wendler-etal-2024-llamas; schut2025multilingual. At each Transformer layer, we project the hidden state $\boldsymbol{h}$ into the vocabulary space using the unembedding matrix $\mathbf{W}^{\text{unembed}}$ and apply a softmax:
$\mathrm{LogitLens}(\boldsymbol{h})=\mathrm{softmax}\left(\mathbf{W}^{\text{unembed}}\,\boldsymbol{h}\right).$
The resulting distribution gives the probability of every token at that layer nostalgebraist2020logitlens. For a specific diagnosis‑label token $t$ (e.g., Protanopia or Deuteranopia) we record
$P(t)=\frac{\exp\left(\boldsymbol{W}_{t}^{\text{unembed}}\cdot\boldsymbol{h}\right)}{\sum_{j}\exp\left(\boldsymbol{W}_{j}^{\text{unembed}}\cdot\boldsymbol{h}\right)}.$
Plotting $P(t)$ across layers reveals how strongly the model favors each color‑vision diagnosis while processing the plate, offering insights into its internal decision-making process.
3.5 Models
We evaluate six LVLMs: Phi-3.5 abdin2024phi3technicalreporthighly, Qwen2.5-VL bai2025qwen25vltechnicalreport, mPLUG-Owl3 ye2024mplugowl3longimagesequenceunderstanding, LLaVA-NeXT liu2024llavanext, Llama-3.2 grattafiori2024llama3herdmodels, and GPT-4o openai2024gpt4ocard. GPT-4o is proprietary; the remaining five are open-source and were selected for passing the Pilot Study prerequisite. Appendix A provides the detailed settings.
Normal Gold 12 8 6 29 57 5 3 15 74 2 6 97 45 5 7 16 73 N/A N/A N/A N/A 26 42 35 96
Llama 19 8 8 42 7 1 2 4 15 2 42 42 1 7 7 4 53 15 15 3 4 52 52 52 8
LLaVA 12 8 6 29 37 3 6 16 7 2 6 9 5 6 7 16 3 1 1 1 1 28 42 35 96
mPLUG 12 8 6 29 37 5 3 15 24 2 6 9 10 5 7 16 13 100 100 10 100 23 42 35 96
Phi
Qwen 12 8 6 29 57 74 74 74 74 74 74 74 74 74 74 74 74 74 74 74 74 74 48 74 95
GPT 12 8 6 29 57 5 3 15 74 2 6 74 45 5 7 16 73 74 74 74 74 5 42 35 96
Llama 19 2 5 17 3 3 2 6 3 2 2 2 45 7 4 6 46 73 4 53 52 4 4 4 2
LLaVA 12 8 6 29 57 3 6 16 7 2 6 9 1 6 7 16 3 1 1 1 1 20 42 38 96
mPLUG 12 8 6 29 37 5 3 15 24 9 6 9 10 5 7 16 13 1 1 1 1 23 42 35 96
Phi
Qwen 21 8 9 29 74 74 74 74 74 74 74 74 74 74 74 74 7 74 74 74 74 38 43 74 95
GPT 12 3 3 29 55 3 5 15 4 9 3 37 45 N/A 7 16 13 74 2 74 27 25 N/A 39 96
Llama 19 2 5 52 1 53 52 49 45 53 52 52 36 55 7 56 4 4 2 49 53 42 52 37 8
LLaVA 12 8 6 29 57 3 3 16 14 2 6 9 10 6 7 16 3 10 1 1 12 20 42 38 96
mPLUG 12 8 6 29 37 5 3 15 24 2 6 9 10 5 7 16 13 1 1 10 1 23 42 35 96
Phi
Qwen 21 8 9 29 74 74 74 74 74 74 74 74 74 74 74 53 7 74 74 74 74 88 43 30 95
GPT 12 8 6 29 57 3 8 N/A 74 2 6 97 45 3 7 16 73 74 N/A N/A 5 93 42 35 96
Llama 19 2 5 17 1 3 2 4 1 3 4 42 46 4 7 3 8 4 4 4 1 3 4 3 8
LLaVA 12 8 6 29 37 3 3 16 7 2 6 9 1 6 7 16 3 1 1 1 12 28 42 38 96
mPLUG 12 8 6 29 37 5 3 15 24 9 6 9 10 5 7 16 13 10 10 10 10 23 42 35 96
Phi
Qwen 21 8 6 29 37 74 74 74 74 74 74 74 74 5 3 53 3 74 0 74 74 38 43 30 95
GPT 12 8 6 29 57 N/A 3 15 74 2 3 37 36 5 7 16 33 74 74 6 3 N/A 42 35 96
Table 5: Predicted outputs of six LVLMs on Ishihara color-vision test plates under four simulated conditions (Normal, Protanopia, Deuteranopia, and Tritanopia) using the Base prompt only. The “Gold” row shows the ground-truth digits. Plates marked as “N/A” indicate those without a correct digit (e.g., blank or pattern-only plates) and are excluded from accuracy calculations.
4 Results and Analysis